### v1

# Multi-Agent Retrieval-Augmented Framework for Evidence Based Counterspeech Against Health Misinformation

Abstract
Large language models (LLMs) incorporated with Retrieval-Augmented Generation (RAG) have demonstrated powerful capabilities in generating counterspeech against misinformation. However, current studies rely on limited evidence and offer less control over final outputs. To address these challenges, we propose a Multi-agent Retrieval-Augmented Framework to generate counterspeech against health misinformation, incorporating multiple LLMs to optimize knowledge retrieval, evidence enhancement, and response refinement. Our approach integrates both static and dynamic evidence, ensuring that the generated counterspeech is relevant, well-grounded, and up-to-date. Our method outperforms baseline approaches in politeness, relevance, informativeness, and factual accuracy, demonstrating its effectiveness in generating high-quality counterspeech. To further validate our approach, we conduct ablation studies to verify the necessity of each component in our framework. Furthermore, human evaluations reveal that refinement significantly enhances counterspeech quality and obtains human preference.
1 Introduction
The rapid spread of misinformation on social media has become a significant concern, particularly in health-related topics. Health misinformation about disease prevention and untested treatments spreads widely, confusing the public and undermining trust in medical institutions (Chou et al., 2021; Cinelli et al., 2020). Consequently, individuals hesitate to follow public health guidelines and some even resort to harmful self-medication practices(Pennycook et al., 2020), underscoring the emergence of mitigating health misinformation. Recent studies prove that social media users can help with misinformation debunking (Chuai et al., 2024; Caulfield et al., 2020), who can rapidly respond, share factual information and correct misinformation, extending the impact of debunking efforts beyond official channels. However, they often struggle to access reliable evidence and are limited by the large volume and complexity of health misinformation (Mourali & Drake, 2022). Therefore, empowering users to effectively counter misinformation is crucial (He et al., 2023; Swire-Thompson et al., 2020).
Large language models (LLMs) are powerful in response generation and are widely used in mitigating misinformation (Proma et al., 2025; Chen & Shu, 2024; Saha et al., 2024; Hong et al., 2024). However, LLMs are not always reliable and can produce hallucinations, prompting researchers to integrate external knowledge to counter misinformation(Bondarenko & Viehweger, 2024). Retrieval-augmented generation (RAG) has been employed to incorporate external knowledge, facilitating the generation of evidence-based counterspeech(Ni et al., 2025; Yue et al., 2024; Liu et al., 2024). Various studies typically rely on static knowledge (Chung et al., 2021; Guu et al., 2020), which refers to verified knowledge that is relatively stable and does not frequently change(Shrinivasan & Razniewski, 2022). For example, the U.S. Surgeon General’s Advisory on Building a Healthy Information Environment  [^1] https://www.ncbi.nlm.nih.gov presents established public health principles and guidelines for addressing health misinformation. Recently, dynamic knowledge has been increasingly incorporated into response generation(Yeginbergen et al., 2025; Komeili et al., 2021; Nakano et al., 2021). This kind of knowledge usually refers to information that is continually updated and responsive to evolving situations. For instance, the U.S. Department of Health and Human Services (HHS) has announced that it will produce 4.8 million doses of the H5N1 avian flu vaccine to prepare for a possible pandemic. This development has been reported by reliable sources such as CIDRAP [^2] https://www.cidrap.umn.edu/, CBS News [^3] https://www.cbsnews.com/, and AJMC [^4] https://www.ajmc.com/. Although recent studies have incorporated various forms of knowledge, most existing methods utilize only one type of knowledge: either the static or the dynamic, ignoring the potential benefits of interacting with both.
Incorporating both static and dynamic knowledge is essential in mitigating health misinformation. Firstly, they are complementary to each other: Although the static knowledge is reliable, it has the limitation that it can not quickly adapt to emerging health misinformation or recent updates. In contrast, dynamic evidence can adapt rapidly to new contexts, but may lack nuance or rigorous validation, potentially weakening the effectiveness of counterspeech(Braun et al., 2024). Therefore, integrating both forms of knowledge allows static knowledge, such as established guidelines, to help filter out low-quality information, while dynamic knowledge ensures the latest development(Charlotte et al., 2024; Nezafat & Samet, 2024). Secondly, incorporating both types of evidence helps build public trust. Individuals are more likely to follow advice relying on longstanding, peer-reviewed research while simultaneously staying alert to contemporary changes(Yeginbergen et al., 2025; Drolsbach et al., 2024).
In our study, we design a multi-source knowledge framework that encompasses both static and dynamic evidence. The static knowledge base provides reliable, focused information, built through careful manual curation. Meanwhile, the dynamic knowledge base continuously integrates real-time updates using a web search API, ensuring that the framework remains responsive to evolving health information. However, incorporating large volumes of evidence may lead to information overload, resulting in responses less accessible to users. To address this, we design a refined module to post-edit responses, ensuring they are clear, concise, and user-friendly(Madaan et al., 2023; Deng et al., 2023). In our study, we propose a Multi-Agent Retrieval-Augmented Framework that integrates various components to achieve effective counterspeech generation. Multiple agents are designed to perform various tasks, including retrieving evidence from static and dynamic knowledge bases, summarizing evidence, generating responses, and refining them for clarity and accessibility.
Our contributions in this study include:
  • Multi-Agent Retrieval-Augmented Framework
    We propose a multi-agent pipeline that integrates a designed retrieval and generation module to realize effective counterspeech generation. Our method outperforms across all evaluation dimensions (Politeness (0.88), and Informativeness (0.78) while maintaining strong Relevance (0.70) and Factual Accuracy (0.86)).
  • Evidence Enhancement and Refined Response Generation
    Our system integrates static retrieval and dynamic retrieval to gather reliable and updated evidence, ensuring robust support for countering health misinformation. Additionally, our refinement agent makes counterspeech more accessible, leading to higher human preference in our evaluation.
2 Related Work
2.1 Evidence-based Misinformation Mitigation
Recently, several studies have leveraged LLMs to produce contextually relevant and effective responses by utilizing their own knowledge base (Mou et al., 2024; Lodovico Molina et al., 2024). However, due to the limitation of LLMs’ knowledge base, they may produce hallucinations and fabricated information (Huang et al., 2025; Rawte et al., 2023). Retrieval-augmented generation (RAG) is proposed to introduce external evidence for producing counterspeech (Ni et al., 2025; Yue et al., 2024; Leekha et al., 2024), improving the factual accuracy of the content generated. This method primarily depends on its knowledge base to provide factual information, but it does not account for health-related changes over time (Ouyang et al., 2025). Some studies combine an LLM with online web searches to identify evidence-based misinformation (Tian et al., 2024; Li et al., 2024a; Dey et al., 2024), which will mitigate the limitations of static knowledge base. These systems enhance evidence retrieval by incorporating information from diverse online sources (Shi et al., 2025). However, web searches introduce new challenges like retrieving low-quality or misleading sources. To address these issues, our study integrates both static and dynamic retrieved evidence to ensure the availability of relevant, real-time information for countering misinformation.
2.2 LLM Agents for Social Response Generation
Integrating LLMs across multiple agents has dramatically improved response generation (Zhang et al., ; Zhao et al., 2024). Multi-agent LLM systems feature several specialized agents working together to tackle complex tasks, utilizing LLMs to boost the quality and efficiency of interactions (Yang et al., 2025; Guo et al., 2024; Cruz, 2024). Many Studies utilize multiple LLM-based multi-agent architectures to improve information retrieval and response generation (Salve et al., 2024), tackling tasks that require long-context understanding (Zhang et al., ), and enhancing recommendation quality and user engagement (Fang et al., 2024). For example, Clarke et al. (2022) implemented a framework where agents specialize in different tasks but coordinate to deliver cohesive responses, utilizing a central management system. These multi-agent approaches demonstrate the effectiveness of breaking down complex response generation into modular, specialized components. Inspired by this, our Multi-agent Retrieval-Augmented Framework employs sequentially organized agents, each with distinct capabilities, to retrieve, filter, summarize, generate, and refine evidence-based counterspeech.
Figure 1: Overview of our proposed Multi-Agent Counterspeech Generation Framework for addressing health misinformation. First, the Static Retrieval Agent gathers information from a local database, while the Dynamic Retrieval Agent fetches real-time evidence online. Next, the Summarization Agent filters and condenses the information for clarity. The Counterspeech Agent then creates a response, and the Refinement Agent improves to ensure it is clear, respectful, and well-supported. This process ensures effective, evidence-based counterspeech.
3 Multi-Agent Retrieval-Augmented Framework
3.1 Generation
The counterspeech generation process involves leveraging LLMs to create effective, informative, and polite responses that counter health misinformation. The primary objective is to generate responses that not only provide factual corrections but also address misinformation in a respectful and engaging manner. Our proposed multi-agent system uses specialized agents working in sequence to retrieve, filter, and process evidence, optimizing the RAG pipeline.
Information Retrieval To incorporate static and dynamic evidence, two separate agents are employed to retrieve evidence: one is a Static Retrieval Agent and another one is a Dynamic Retrieval Agent. These Agents are collecting information based on the misinformation posts. The Static Retrieval Agent conducts an inquiry within a structured local knowledge base by employing a dual approach that integrates vector-based semantic search with keyword-based exact matching (Yuan et al., 2024; Sawarkar et al., 2024), ensuring that only valid medical knowledge will be included in this process. In parallel, the Dynamic Retrieval Agent performs real-time web searches using the DuckDuckGo API, introducing the latest health misinformation corrections from authoritative sources. Two sources of knowledge are combined into the final evidence.
Evidence Summarization To prevent information overload from combined evidence, we introduce a Summarization Agent to filter and distill the most relevant details combined. This step is crucial for eliminating redundancy and resolving conflicting information, ensuring that only the most reliable and pertinent content is retained for counterspeech generation. Specifically, we prompt the Summarization Agent to condense evidence into a clear and refined summary.
Counterspeech Generation We integrate distilled evidence from the Summarized Agent into generation by prompting Counterspeech Generation Agent to produce evidence-based counterspeech.
Counterspeech Refinement To further enhance accessibility and clarity, we employ the Refinement Agent to reframe and polish counterspeech, ensuring that the generated counterspeech is not only well-supported by evidence but also contextually relevant and trustworthy.
3.2 Evaluation
We employed a comprehensive evaluation framework based on four key criteria: politeness, relevance, informativeness, and factual accuracy. These criteria ensure that the responses are engaging, respectful, factually sound, and effective in addressing health misinformation.
(a) Politeness: Politeness in responses is essential in communication, which helps to foster user engagement and build a supportive health community(Shan et al., 2022). A polite counterspeech helps avoid potential backlash and is more likely to be accepted by users (Perez et al., 2025), as it encourages a respectful tone(Song et al., 2025a; Yue et al., 2024; He et al., 2023). To evaluate politeness, we utilize the Multilingual Politeness Classification Model (Srinivasan & Choi, 2022) to assess whether a response adheres to polite discourse norms.
(b) Relevance: Relevant responses ensure the response directly addresses the specific misinformation claim (Yue et al., 2024; He et al., 2023). We assess relevance using the BM25 relevance scoring algorithm (Robertson et al., 2009), which measures how well the generated response aligns with the original misinformation content.
(c) Informativeness: Informativeness reflects the amount of useful, relevant, and evidence-based content shared in the response. A highly informative counterspeech response not only corrects misinformation but also provides actionable insights and explanations that improve user understanding and support informed decision-making(Chung et al., 2024). To measure informativeness, we employ a LLM as an evaluator (Li et al., 2024b) with detailed instructions, assessing if the response provides valuable information.
(d) Factual Accuracy: Factual accuracy demonstrates the reliability and trustworthiness of the generated response by ensuring that the information is correct and supported by evidence(Zhou et al., 2024). In counterspeech Generation, presenting scientifically accurate facts can effectively correct misinformation and maintaining user trust (Yue et al., 2024). We employ LLM evaluation (Li et al., 2024b) to check how accurate the generated answers are by comparing them to the gathered evidence. The model rates how accurate a response is by checking if it correctly presents scientific facts and does not include any false information.
4 EXPERIMENTS AND RESULTS
4.1 Dataset
We utilize PRW API  [^5] https://praw.readthedocs.io/ to collect health misinformation from Reddit, focusing on topics related to the coronavirus disease (COVID-19), influenza (Flu), and human immunodeficiency virus (HIV), as these topics have been the center of conspiracy theories and misinformation in recent times. We collect Reddit posts and comments containing health-related keywords (e.g., ”vaccines,” ”COVID-19,” ”alternative medicine”, etc. See details in Appendix A.1.2)). Finally, we obtain about 4,968 posts and 17,223 comments from high-engagement subreddits (e.g., r/health or r/science, etc. See details in Appendix A.1.1).
After collecting the Reddit dataset, we employ human annotators to identify and filter health misinformation (Song et al., 2025b). Considering the annotation cost, we only extract 1000 samples for annotation. Five annotators with backgrounds in information science were recruited for annotating. To ensure consistency and reliability in annotations, all annotators were provided with comprehensive and detailed annotation guidelines Appendix A.2.1. We finally obtained 330 posts containing health misinformation.
To further enrich our dataset, we developed an automatic classifier trained on the annotated data, enabling the efficient collection of additional health misinformation posts from unannotated data. We employ the RoBERTa-large model (Liu et al., 2019), trained on our manually annotated dataset, which achieves a moderate accuracy with an F1 score of 0.76. We apply it to the remaining dataset, identifying 831 additional posts containing health misinformation. To validate the reliability of the classifier, we extract 100 samples to conduct human validation (See details in Appendix A.2.2). The final dataset contained 1,161 posts labeled as health misinformation.
4.2 Baseline
We compared four different baseline methods to determine the most effective way to counter health misinformation against our proposed Multi-Agent counterspeech Generation System.
LLM With Direct Prompt (LLM w DP) This method uses direct instructions (Appendix  A.4.1) to prompt LLM to generate counterspeech without any external knowledge. The model takes health misinformation posts as input and follows direct instructions to generate responses using LLM’s internal knowledge.
LLM With Prompt Engineering (LLM w PE) This method uses the refined Guided Instruction prompt (Appendix A.4.2) to structure the LLM’s responses, ensuring clarity, factual accuracy, and coherence, however, still relying on the LLM’s internal knowledge.
Static RAG This approach employs RAG to extract supporting evidence from our curated knowledge bases, enabling us to validate the effectiveness of the static evidence in generating accurate and evidence-based responses. The retrieved evidence is then integrated into a Guided Instruction prompt (Appendix A.4.2), guiding the LLM in generating evidence-based counterspeech.
Dynamic RAG This approach retrieves real-time evidence from the DuckDuckGo search engine, using the misinformation post as a query. By leveraging real-time web search, this method enables dynamic retrieval of the most recent and relevant information, allowing for a comparative analysis between dynamic evidence and static knowledge bases. Notably, this retrieval process follows the same Guided Instruction prompt framework, allowing a comparative assessment of the effectiveness of dynamic evidence in generating counterspeech.
Figure 2: Effect of different top-k evidence filtering on Multi-Agent performance across various metrics (Polite: politeness, Rel: relevance, Info: informativeness, Fact: factual Accuracy).
4.3 Multi-Agent Retrieval-Augmented Framework Implementation
The Multi-agent Retrieval-Augmented Framework consists of multiple specialized agents that operate sequentially to optimize the retrieval, filtering, summarization, generation, and refinement of responses to health misinformation. We use the ”Llama-3.1-8B-Instruct” (Grattafiori et al., 2024) model for generation. Generation configurations are detailed in Appendix A.3.
Knowledge Source In terms of the static knowledge source, we collect evidence from trusted sources such as the National Institutes of Health (NIH)  [^6] https://www.ncbi.nlm.nih.gov  [^7] https://covid19community.nih.gov World Health Organization (WHO)  [^8] https://www.who.int  [^9] https://www.who.int/docs/default-source  [^10] https://www.who.int/emergencies/diseases. These institutions are globally acknowledged for their authoritative presence and credibility, offering regular updates on health-related information that are crucial for creating trustworthy counterspeech (Kington et al., 2021). Our dynamic knowledge source utilizes the DuckDuckGo search engine API to retrieve real-time, web-based information. This method enables the system to access current evidence from credible news outlets, government health websites, and fact-checking organizations.
Static Retrieval Agent This Agent performed a Hybrid search in our curated knowledge base using both Vector-Based Retrieval and Keyword-Based Retrieval. Vector-based retrieval employs a pre-trained embedding model ”sentence-transformers/all-mpnet-base-v2”  [^11] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for embedding-based similarity searches and ranking by cosine similarity, while Keyword-Based Retrieval utilizes ”SimpleKeywordTableIndex” to identify exact team matches within the documents.
Dynamic Retrieval Agent The Dynamic Retrieval Agent uses the DuckDuckGo API [^12] https://pypi.org/project/duckduckgo-search-api/ to fetch live external evidence. The misinformation post serves as the structured search query. It extracts key factual sentences from each result via web scraping (using BeautifulSoup [^13] https://pypi.org/project/beautifulsoup4/). By querying DuckDuckGo, the Dynamic Retrieval Agent can capture the latest posts, news articles, and trusted websites relevant to a given health misinformation post. Additionally, we conduct fact-checking and filter the information with lower factual accuracy, ensuring that the selected information maintains an average accuracy score above 0.65.
Summarization Agent This Agent ranks the retrieved evidence collected from the Static Retrieval Agent and the Dynamic Retrieval Agent. We evaluate various top-k evidence selections, as illustrated in Figure 2, which indicates that top-10 evidence yields a slightly better performance across all dimensions. Therefore, we filter the top-10 evidence using BM25-based ranking and semantic embedding similarity and summarize them using the prompt in Appendix  A.4.3, effectively preserving key factual details.
Counterspeech Generation Agent The summarized evidence is then further utilized by the LLM to generate fact-based counterspeech. This agent employs a structured prompting approach (Appendix  A.4.2), which effectively guides the LLM to produce concise, informative, and contextually relevant responses.
Refinement Agent The final step involves refining the generated responses using the Refinement Agent. A refined prompt (Appendix A.4.4) guides this process, ensuring that responses are respectful, engaging, and persuasive.
Method Politeness Relevance Informativeness Factual Accuracy
LLM w DP 0.44 (0.26) 0.70 (0.14) 0.77 (0.11) 0.81 (0.21)
LLM w PE 0.84 (0.15) 0.65 (0.15) 0.75 (0.07) 0.83 (0.20)
Static RAG 0.80 (0.15) 0.65 (0.16) 0.77 (0.08) 0.81 (0.24)
Dynamic RAG 0.86 (0.16) 0.64 (0.15) 0.73 (0.11) 0.83 (0.19)
Multi-Agent (Ours) 0.88 (0.14) 0.70 (0.13) 0.78 (0.13) 0.86 (0.19)
Table 1: Comparison of Multi-Agent counterspeech generation with baseline methods across key evaluation metrics. (DP: direct prompt, and PE: prompt engineering.) Higher values indicate better performance. The best scores in each category are underlined.
4.4 Evaluation Results
Table 1 presents the results of the experiment. The overall performance analysis indicates that LLM w DP exhibits slightly strong relevance (0.70) and informativeness (0.77), compared with other baseline methods. Prompt engineering (LLM w PE) significantly improves politeness (0.84) and factual accuracy (0.83) compared with LLM w DP. However, it comes at the cost of slightly lower relevance (0.65).
Introducing static knowledge retrieval (Static RAG) provides better grounding by integrating factual information from a structured knowledge base. However, the results indicate that this does not result in a significant overall improvement across all dimensions compared with LLM without external knowledge (LLM w DP and LLM w PE), suggesting that LLM may struggle to effectively incorporate external knowledge into generation (Bondarenko & Viehweger, 2024). This pattern remains consistent in the Dynamic RAG experiment (see Figure 3), even when real-time external evidence is provided, further validating that LLMs may have inherent limitations in effectively integrating external knowledge.
Our Multi-agent Retrieval-Augmented Framework successfully addresses these gaps by effectively integrating multiple retrieved evidence and refining the final responses. As a result, it achieves higher politeness (0.88), informativeness (0.78), and factual accuracy (0.86) while maintaining substantial relevance (0.70), making it a more effective and balanced approach to countering misinformation.
Figure 3: Performance of counterspeech generation approaches across evaluation metrics.
4.5 Ablation Experiments
To demonstrate the necessity of multiple agents in our framework, we conduct ablation experiments to generate counterspeech. Additionally, we explore various prompting strategies to assess whether they lead to performance improvements. Table 2 represents how the involvement of different agents contributes to the performance of counterspeech generation under the Guided and CoT prompting strategies (See in Appendix A.4).
MA w/o SA w/o RF shows the effect after the removal of Summarization and Refinement Agents. This setup only relies on retrieval and response generation without any additional refinement. It maintains relatively high scores for factual accuracy (Guided: 0.85) and informativeness (Guided: 0.78), however, with lower politeness (Guided: 0.80) compared with MA. Including the Summarization Agent (MA w/o RF with Guided) improves the relevance (0.70), factual accuracy (0.87), and informativeness (0.82) compared with MA w/o SA w/o RF. This further demonstrates that summarized evidence helps LLMs better integrate evidence into generation. Furthermore, the addition of the Refinement Agent, our proposed Multi-Agent approach (MA (Ours)), further refines responses by ensuring they are organized, polite, and clear, with the best performance across all dimensions.
We also examine the effect of changing the Prompting Strategy to Chain-of-Thought (CoT) (See details in Appendix A.4.5). While the CoT prompting strategy enables the generation to maintain similar factual accuracy compared with Guided prompting, the performance of politeness (MA w/o SA w/o RF: 0.63 vs 0.80; MA w/o RF: 0.65 vs 0.79; MA: 0.74 vs 0.88), and informativeness (MA w/o SA w/o RF: 0.76 vs 0.78; MA w/o RF: 0.77 vs 0.82; MA: 0.77 vs 0.78) decrease. The results indicate that CoT reasoning may be less effective than Guided prompting in our study. The step-by-step reasoning process often introduces redundant or overly detailed content, which decreases politeness and informativeness (Sprague et al., 2024; Wei et al., 2022).
Method Prompt Politeness Relevance Informativeness Factual Accuracy
MA w/o SA w/o RF CoT 0.63 (0.24) 0.69 (0.13) 0.76 (0.17) 0.84 (0.18)
MA w/o SA w/o RF Guided 0.80 (0.19) 0.67 (0.19) 0.78 (0.13) 0.85 (0.19)
MA w/o RF CoT 0.65 (0.22) 0.69 (0.13) 0.77 (0.09) 0.86 (0.18)
MA w/o RF Guided 0.79 (0.17) 0.70 (0.13) 0.82 (0.15) 0.87 (0.19)
MA CoT 0.74 (0.19) 0.67 (0.14) 0.77 (0.09) 0.87 (0.19)
MA (Ours) Guided 0.88 (0.14) 0.70 (0.13) 0.78 (0.13) 0.86 (0.19)
Table 2: Evaluation of the impact of Summarization Agent (SA) and Refinement Agent (RF) alongside different prompting styles (Guided vs. CoT) on Multi-Agent (MA) counterspeech generation performance.
5 Human Evaluation
We conduct two rounds of human evaluations. The first round is to validate the LLM evaluation. Even though various studies rely on LLMs as evaluators(Li et al., 2024b; Evans et al., 2024; Chen et al., 2024), different research tasks can yield varying LLM performance (Chern et al., 2024; Chan et al., 2023). Therefore, we involve humans to verify LLM evaluation of informativeness and factual accuracy. The second round aims to explore the effectiveness of the refined response. In our study, we utilize the Refinement agent to improve counterspeech, however, it remains unknown whether users prefer these refined responses. Therefore, it is necessary to figure out human preference, which may help us improve generation in future studies.
In the first stage, we conduct a human evaluation to verify the reliability of the computation evaluation: Informativeness and Factual Accuracy. The primary objective is to determine how well the model-generated evaluation scores align with human judgments. We randomly select 100 samples, and three annotators independently evaluate the generated responses based on informativeness and factual accuracy on a 1-5 scale (Details guidelines in Appendix A.5). Initially, the broad rating scale results in variability in annotator agreement. To address this, further discussions among the annotators are held, leading to a final label that reflects a consensus-driven evaluation. Once the final human labels are established, we compared them with our model evaluation results. For factual accuracy, Cohen’s Kappa is $\kappa\geq 0.69$, and for informativeness, Cohen’s Kappa is $\kappa\geq 0.65$, indicating substantial agreement with the model evaluation. These results suggest that the computing-based evaluation exhibits strong consistency with human judgments.
Secondly, we assess the effectiveness of the Refinement Agent in counterspeech responses. We conduct a comparative human evaluation of two different response strategies: Response A (Multi-Agent w/o Refinement Agent) and Response B (Multi-Agent w Refinement Agent). The goal is to determine human preference and assess how the Refinement Agent enhances response quality(details of the guidelines in Appendix A.5). The initial agreement rates among them are 74%, 71%, and 73%, respectively, which shows they have a relatively high level of consistency. The disagreements were solved by further discussion and concluded with a final label. The annotation results reveal that 72% of users prefer Response B, which deploys a Refinement Agent, indicating that clearer and better-organized responses are more accessible for users. In the future, we can continue refining our methods to produce counterspeech that is even more user-friendly and effective.
6 CONCLUSION
In this study, we present a Multi-Agent Retrieval-Augmented Framework for generating effective counterspeech against health misinformation. This framework integrates multiple specialized LLM agents to sequentially (1) retrieve evidence from both static knowledge bases and dynamic web sources, (2) filter and summarize the combined evidence, and (3) generate and refine counterspeech using structured prompting. This design enables the generation of responses that are not only factually accurate but also polite, informative, and contextually relevant. Experimental results and human evaluations demonstrate that our method outperforms several baselines across multiple dimensions. Additionally, ablation studies confirm the importance of each module, particularly the refinement agent, which plays a key role in enhancing the clarity and engagement of counterspeech. Our findings highlight the value of combining static and dynamic knowledge and the refined method to improve the reliability and quality of automated counterspeech. We acknowledge several limitations in our study. Even though we incorporate both static and dynamic evidence, we recognize that we have not fully collected all existing static knowledge since the cost of manual work. Moreover, there may be some overlap between the two knowledge sources. Nevertheless, we claim that each knowledge base serves a distinct purpose: static evidence offers rigorously validated information, while dynamic evidence provides real-time updates. And the proposed Multi-Agent Retrieval-Augmented Framework enables efficient integration of these components to generate effective counterspeech. In future work, we aim to further refine the curation processes for both knowledge sources and explore methods to minimize overlap, enhancing the accuracy of our framework in addressing health misinformation.
Ethics Statement
Our study follows the highest ethical standards to address misinformation using Large language models. We acknowledge that AI-generated responses can affect society, and we are careful to ensure they are effective while still being ethical. We use publicly available data sources, ensuring that no personally identifiable information is collected or used. User identities from social media datasets are anonymized, and any sensitive information is removed before analysis. Human annotators involved in data annotation and evaluating counterspeech responses remain anonymous, and their participation is voluntary. We follow rules to make sure the annotation process is fair and open. Our study aims to generate respectful, fact-based counterspeech while minimizing unintended biases. We understand that LLMs can show biases based on the data on which they were trained. To address this, we use retrieval-augmented methods to ensure our answers are based on verified facts. We ensure that the AI-generated responses are not misleading, confrontational, or harmful. The system is created to give polite and well-structured responses that encourage helpful discussion rather than increasing division. We are committed to transparency in our methodology, which includes providing details on model architectures, data sources, and evaluation criteria. Our work aligns with the principles of ethical AI research and aims to positively contribute to public discourse by countering misinformation with reliable evidence. By following these ethical considerations, We aim to create AI solutions that encourage good communication, build public trust, and deliver accurate, respectful, and reliable information.
Appendix A Appendix
A.1 Subreddits and Keyword lists
A.1.1 List of Subreddits used
Table 3 presents the list of subreddits analyzed in this study. These subreddits were selected to cover a broad spectrum of discussions on science, public health, medical topics, vaccines, and conspiracy theories.
Subreddit Name
r/science, r/Wuhan_Flu, r/CoronavirusCirclejerk, r/DebateVaccines, r/vaccinelonghaulers,
r/ChurchofCOVID, r/vaccineautismevidence, r/ScienceUncensored, r/CoronavirusUS, r/conspiracy,
r/medical_advice, r/joerogan, r/Conservative, r/vaccineaddiction, r/thingsprovaxxerssay
Table 3: List of subreddits analyzed in the study.
A.1.2 List of keywords used
Table 4 presents the list of keywords used to extract posts and submissions from the selected subreddits. These keywords encompass a wide range of medical, health, and pandemic-related discussions.
Keywords
diagnosis, treatment, prescription, prognosis, nutrition, diet, exercise, chemicals, abortion, vaccine, jabs,
COVID-19, surgery, pills, symptom, immunity, disease, autism, cure, infection, infected, pandemic, health,
Corona, virus, medical, doctor, insomnia, side effects, polio, symptoms, fibromyalgia, acidity, cancer,
tuberculosis, HIV, medication, contraception, immunocompromised, covid, vax, truth, sick, flu, pregnancy,
pneumonia, measles, jab.
Table 4: List of keywords used to extract relevant posts from subreddits.
A.2 Guidelines and Annotation process for detecting health misinformation
A.2.1 Annotation guidelines and process
After collecting the Reddit dataset, the next step was to annotate posts into two categories: ”Health Misinformation” and ”Not Health Misinformation.”. To ensure annotation consistency, all annotators received comprehensive guidelines, including examples of health misinformation, classification criteria, and citations of reputable sources (e.g., CDC, WHO, and NIH) for verification purposes.
Annotation guidelines for health misinformation labeling
The goal of this annotation task is to classify posts based on whether they
contain health-related misinformation. Posts will be assigned one of two labels:
Label 1: "Health Misinformation", If the post contains health-related
misinformation.
Label 0: "Not Health Misinformation", If the post does not contain any
health-related misinformation or is unrelated to health information.
Definition of Health Misinformation:
Any false, misleading, or unverified claim related to health, medicine,
diseases, treatments, vaccines, nutrition, or wellness.
Misinformation includes claims that contradict established medical research,
public health guidelines, or authoritative sources such as the World Health
Organization (WHO) and the Centers for Disease Control and Prevention (CDC).
Annotation process
The annotation process was conducted in two stages to ensure consistency and accuracy. In the first stage, three annotators independently labeled 480 posts as either health misinformation or not health misinformation. In the second stage, another set of three annotators annotated 499 additional posts using the same classification scheme. Each annotator independently reviewed the content and fact-checked it against authoritative sources before assigning labels.
Despite the structured guidelines, some posts posed challenges in classification, particularly those containing satirical statements, ambiguous claims, or partially correct yet misleading information. In the first stage, the Cohen’s Kappa agreement rates between annotators were $\kappa\geq 0.58$, $\kappa\geq 0.67$, and $\kappa\geq 0.54$, indicating moderate to substantial agreement. In the second stage, the agreement rates were $\kappa\geq 0.69$, $\kappa\geq 0.48$, and $\kappa\geq 0.56$, also reflecting a mix of moderate and substantial agreement. To resolve annotation disagreements, the annotators engaged in group discussions and majority voting, ensuring that the final classification was well-supported by evidence. In cases where uncertainties persisted, the posts were cross-checked against fact-checking websites such as Snopes, HealthFeedback, and FactCheck.org to verify the accuracy of the information. This multi-step validation process ensured that the final dataset maintained a high level of accuracy and reliability.
After resolving all disagreements and finalizing the labels, the dataset consisted of 330 posts labeled as health misinformation and 648 posts labeled as not health misinformation.
A.2.2 Human validation
A human validation step was conducted to validate the reliability of these automatically labeled posts. Three annotators independently fact-checked a random sample of 100 model-labeled posts, assessing their accuracy. The agreement rate between the annotators is 88.1%, 89.1%, and 85.1%, and Cohen’s Kappa scores are $\kappa\geq 0.73$, $\kappa\geq 0.75$, and $\kappa\geq 0.67$, respectively. Which shows substantial agreement between the annotators. For the disagreement, we conducted further discussion and fact-checking, which concluded us with the final label. With the final label, our model agreement rate is 88.1% and Cohen’s Kappa, $\kappa\geq 0.73$ demonstrates a substantial agreement between the model’s classification and human judgment.
A.3 Computing resources and model parameters
A.3.1 Hardware configuration
We utilized a multi-GPU setup with the following specifications:
  • CPU: Intel® Xeon® Gold 6226R @ 2.90GHz
    • Cores: 16 per socket, total 32 cores
    • Threads: 64 (2 threads per core)
  • Memory: 125GB RAM
  • GPU: Three NVIDIA Quadro RTX 8000 GPUs
    • CUDA Version: 12.6
    • Memory: 48GB per GPU (Total 144GB VRAM)
Our Multi-agent Retrieval-Augmented Framework was deployed across three GPUs, leveraging parallel processing for retrieval, summarization, and response generation. To ensure efficiency, data pre-processing was optimized through multithreading, and retrieval mechanisms combined keyword-based and embedding-based search strategies.
A.3.2 Model and hyperparameter settings
The key hyperparameter settings for all experiments are as follows:
  • LLM: Meta-Llama 3.1-8B-Instruct [^14] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
  • Embedding Model: Sentence-Transformers (all-mpnet-base-v2) [^15] https://huggingface.co/sentence-transformers/all-mpnet-base-v2
  • Summarization:
    • Temperature: 0.3
    • Top-p: 0.8
  • Counterspeech Generation:
    • Temperature: 0.6
    • Top-p: 0.85
  • Response Refinement:
    • Temperature: 0.4
    • Top-p: 0.85
A.4 Prompt Design
A.4.1 Direct instructions prompt for counterspeech generation:
Generate a counterspeech response to the given health misinformation post:
Post: {post_content}
Counterspeech:
A.4.2 Prompt Engineering (Guided prompt) for counterspeech generation:
Guided prompt:
You are an AI designed to provide concise, respectful, and evidence-based
counterspeech to health misinformation in the form of a comment responding
directly to a post.
Guidelines:
- Begin with an empathetic tone to validate user concerns before addressing
  the misinformation.
- Clearly and succinctly correct misinformation using simple, factual
  explanations.
- Integrate concise evidence from trusted sources (e.g., CDC, WHO, NIH) with
  clear attribution, limiting to 2-3 references.
- Avoid redundant information, unrelated anecdotes, or excessive links.
- Craft responses that are precise, easy to understand, and actionable,
  keeping them under 120 words in a direct reply format.
Counterspeech generation prompt (static & dynamic):
{Guided Prompt}
Post: {post_content}
Evidence: {evidence}
Response:
Counterspeech generation prompt (Multi-Agent):
{Guided Prompt}
Misinformation Post:
{post_content}
Verified Evidence:
{summarized_evidence}
Write a counterspeech response that:
- Directly refutes the misinformation in a confident yet respectful manner.
- Provides an example that explains why the misinformation is incorrect.
counterspeech Response:
A.4.3 Summarization Agent prompt
Below is a collection of factual evidence addressing misinformation:
{combined_filter_evidence}
Summarize this evidence concisely, preserving key factual
details while removing redundant or speculative information.
Final Summarized Evidence:
A.4.4 Refinement Agent prompt
Below is a counterspeech response to a misinformation Reddit post.
Refine the response to ensure it is clear, factually accurate, and
respectful, without reducing its effectiveness in addressing the misinformation.
Misinformation Post:
{post_content}
Current counterspeech Response:
{response}
Refinement Guidelines:
- Begin with an empathetic and respectful tone.
- Be more polite, respectful while clearly addressing the misinformation.
- Improve clarity and readability while keeping key factual evidence intact.
- Ensure the final response is factually accurate without removing or altering
any factual information.
- Add references from trusted sources if available.
- Keep the response concise (under 120 words), but make sure it does not
lose relevance.
Refined Response:
A.4.5 Chain-of-Thought (CoT) prompt for generation
Chain-of-Thought (CoT) counterspeech generation prompt:
You are an AI fact-checking specialist designed to analyze health
misinformation and generate counterspeech through structured reasoning.
Your process must:
1. Maintain scientific accuracy while being accessible
2. Prioritize harm reduction in responses
3. Balance empathy with factual rigor
4. Adhere to medical consensus from WHO/CDC/NIH
Health Misinformation Post:
{post_content}
Verified Evidence:
{summarized_evidence}
Generate a counterspeech response to this Health Misinformation Post
using the following reasoning process:
1. Analyze the post to identify key claims and emotional undertones.
2. Select the most relevant facts from verified evidence.
3. Explain why the claims are misleading using:
   a) Logical contradictions in the claims.
   b) Evidence from authoritative sources.
   c) Concrete examples/counter-examples.
4. Structure the response with:
   - Empathetic acknowledgement.
   - Clear factual correction.
   - Simple analogy/metaphor for understanding.
   - Keep the response concise, under 150 words.
   - Trusted source references.
Reasoning Process:
[Think through each step carefully]
counterspeech Response:
Chain-of-Thought (CoT) prompt for counterspeech refinement:
Refine the generated counterspeech response to a health misinformation post
while maintaining clarity, factual accuracy, and politeness.
1. Identify and resolve logical contradictions.
2. Strengthen arguments with source-attributed evidence.
3. Improve structure for effective communication (acknowledgment  correction
 analogy).
Misinformation Post:
{post_content}
Current counterspeech Response:
{response}
Refinement Guidelines:
- Rewrite with a respectful and empathetic tone.
- Ensure analogies/examples are universally understandable.
- Remove redundant explanations while ensuring clarity.
- Trim to 120 words without losing key evidence.
Output Format:
[Empathetic and Respectful Opening]
[Clear Factual Correction]
[Simple Analogy/Example]
[Trusted Source References]
Refined Response:
A.5 Annotation guidelines for factual accuracy and informativeness assessment
Guideline for informativeness (1-5 Scale) :
Evaluate the informativeness of the following counterspeech response on a scale
of 1 to 5.
Score 1: Not Informative: The response lacks substance or is irrelevant.
Score 2: Slightly Informative: Some relevant points but too vague or missing
key details.
Score 3: Moderately Informative: Provides a basic understanding but lacks
depth.
Score 4: Mostly Informative: Covers most aspects well, with minor gaps.
Score 5: Completely Informative: Comprehensive, well-structured, detailed,
and highly informative.
Factual accuracy evaluation guidelines (1-5 Scale):
Your task is to assess the factual accuracy of the counterspeech response using
the provided evidence. Compare the response against the evidence and determine
its correctness using the scoring criteria below. You need to verify the
link/reference (if available) in the counterspeech response and also need to
write a comment explaining why you chose this score.
Score 1: Completely Inaccurate: Mostly inaccurate, contains false claims.
Score 2: Mostly Inaccurate: Some correct facts, but also contains misleading
information.
Score 3: Partially Accurate: Mostly correct, but lacks sufficient evidence.
Score 4: Mostly Accurate: Accurate, well-supported, minor inconsistencies.
Score 5: Completely Accurate: Fully fact-based, verified by reliable sources.
A.6 Human evaluation guidelines for counterspeech comparison
Your task is to compare two responses (Response A and Response B) that address
a misinformation post.
You will decide which response is better based on your preference.
How to Evaluate:
1. Read the Misinformation Post carefully to understand the claim being made.
2. Compare Response A and Response B.
3. Select the response you may prefer to use in addressing the misinformation.
4. Provide a brief comment explaining why you selected that response.
Note: Response A (Multi-Agent w/o Refinement Agent) and
      Response B (Multi-Agent w Refinement Agent)